The Robinson Company sells sports decals that can be personalized with a player’s name, a team name, and a jersey number for $6 each. Robinson buys the decals from a supplier for $2.60 each and spends an additional $0.70 in variable operating costs per decal. The results of last month’s operations are as follows Sales Revenue: $12000 Cost of Goods Sold: 5200 Gross Profit: 6800 Operating Expenses: 3800 Operating Income: 3000 Contribution Margin per unit = $2.70 What is Crane's monthly breakeven point in units? In dollars?
Answer #1

Selling price per unit = $6

Sales revenue = $12,000

Number of units sold = Sales revenue/Selling price per unit

= 12,000/6

= 2,000

variable cost per unit = Purchase price + Operating expense

= 2.60 + 0.70

= $3.30

Contribution margin per unit = Selling price per unit - variable cost per unit

= 6 - 3.30

= $2.70

Total variable cost = Number of units sold x variable cost per unit

= 2,000 x 3.30

= $6,600

Cost of goods sold = $5,200

Operating expense = $3,800

Total cost = Cost of goods sold + Operating expense

= 5,200 + 3,800

= $9,000

Fixed cost = Total cost - Total variable cost

= 9,000 - 6,600

= $2,400

Break even point (in units) = Fixed cost/Contribution margin per unit

= 2,400/2.70

= 889 units

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 2.70/6

= 45%

Break even point (in dollar) = Fixed cost/Contribution margin ratio

= 2,400/45%

= $5,333

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.